OpenStudy (iwanttogotostanford):

HELP PLEASE ASAP:

OpenStudy (iwanttogotostanford):

@sshayer @steve816 @triciaal

OpenStudy (steve816):

No symmetry I believe.

OpenStudy (iwanttogotostanford):

@steve816

OpenStudy (iwanttogotostanford):

i need help with like 5 more if you please could :-)

OpenStudy (iwanttogotostanford):

@mathstudent55 @Nnesha

OpenStudy (iwanttogotostanford):

@FaiqRaees

OpenStudy (iwanttogotostanford):

someone please i have so many questions

OpenStudy (iwanttogotostanford):

@calculusxy

OpenStudy (iwanttogotostanford):

@sshayer

OpenStudy (iwanttogotostanford):

@sshayer please help ASAP

OpenStudy (sshayer):

if the asymptote is y=pm a/b x and foci are (0,+-a) then eq. of hyperbola is \[\frac{ y^2 }{ a^2 }-\frac{ x^2 }{ b^2 }=1\] here foci is (0,+-4) a=4 \[\frac{ a }{ b }=\frac{ 1 }{ 3 },\frac{ 4 }{ b }=\frac{ 1 }{ 3 },b=14\] now you can write the eq.

OpenStudy (iwanttogotostanford):

oh xo it is B!!!

OpenStudy (iwanttogotostanford):

@sshayer

OpenStudy (sshayer):

what about A?

OpenStudy (sshayer):

correction b=4*3=12

OpenStudy (iwanttogotostanford):

A is right....

OpenStudy (iwanttogotostanford):

@Kevin can you help me with a couple math questions please??

OpenStudy (kevin):

Sorry, I just about to go sleep

OpenStudy (iwanttogotostanford):

just 1x please?

OpenStudy (kevin):

hmm.... tomorrow? xD

OpenStudy (kevin):

let the other help u.. See you

OpenStudy (iwanttogotostanford):

@Will.H @Nnesha

OpenStudy (iwanttogotostanford):

@ajspeller

OpenStudy (iwanttogotostanford):

@triciaal

OpenStudy (triciaal):

strong suggestion calm down and organize in groups by topic open post with one of each kind pay attention to the steps then apply to the others